OpenStudy (sharkfinsoup):

Which of the following expressions is this one equivalent to? (3x^3+15x^2+17x+3)/(x+5) A. x^3-2x+5+7/(x+5) B. 3x-17+88/(x+5) C. 3x^2+17-82/(x+5) D. 3x^2-5x-8-37/(x+5)

alones (alones):

Alright so you know the formula? \[A/B \implies A= BQ + R\] As \(R\) stands for reminder and \(Q\)

alones (alones):

**And \(Q\) as the quotient..

alones (alones):

\[\frac{ (3x^3+15x^2+17x+3) }{ x+5 = 3x²+17 } \] So, as you see the reminder should be \( - 82\)

alones (alones):

And now, \[\frac{ (3x^3+15x^2+17x+3= (3x²+17) }{ (x+5) }\] So your final answer would be: \[\frac{ (3x²+17)-82 }{ (x+5) } \]

alones (alones):

@sharkfinsoup I'm trying to see if this makes sense, and you don't even reply LOL...Does this makes sense? :'3

OpenStudy (sharkfinsoup):

Yes it does thank you so much. Sorry I didn't respond, my computer froze for some reason and I had to restart it. I really appreciate thank you @AloneS

alones (alones):

Awh haha wow, I froze for a second, and it went to usual. I was glad to help<3

OpenStudy (sharkfinsoup):

:)